OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. M(0, 6), N(6, 0)

OpenStudy (john_es):

Well, the standard form (or implicit form) is, \[A x+ By = C\] But in order to get it, I prefer to obtain this, \[y=mx+n\] First, we find m, the slope, \[m=\frac{0-6}{6-0}=-1\] Then, we have, \[y=-x+n\] We use now one point of the problem, for example, (0,6), then, \[6=-0+n\Rightarrow n=6\] So, you have, \[y=-x+6\] Reordering, \[x+y=6\]